A 8.21 kg bowling ball traveling at 14.66 m/s strikes a 2.1 kg bowling pin. The impact slows the bowling ball to 9.97 m/s.



Determine the change in momentum of the pin.
1
Expert's answer
2020-04-15T10:12:27-0400

"m_0v_0=m_0v+mu"


"\Delta p=mu=m_0v_0-m_0v=m_0(v_0-v)=8.21(14.66-9.97)=38.5\frac{kg\cdot m}{s}"